35 Determine the inverse Laplace transform of: 1 2 32-8
Answer #1

Key points:

  • L^{-1}\left \{ \frac{1}{s-a} \right \}=e^{at}
  • S (-1 = cosh(at) s2 - a2

We have,

L^{-1}\left \{ \frac{3s}{\frac{1}{2}s^2-8} \right \}=L^{-1}\left \{ \frac{6s}{s^2-16} \right \}

=L^{-1}\left \{ \frac{6s}{(s-4)(s+4)} \right \}

=6L^{-1}\left \{ \frac{s}{(s-4)(s+4)} \right \} [Using linearity of inverse laplace transform]

=\frac{6}{2}L^{-1}\left \{ \frac{1}{s-4}+\frac{1}{s+4} \right \}

=3L^{-1}\left \{ \frac{1}{s-4}+\frac{1}{s+4} \right \}

=3L^{-1}\left \{ \frac{1}{s-4}\right \}+3L^{-1}\left \{\frac{1}{s+4} \right \}  [Using linearity of inverse laplace transform]

=3e^{4t}+3e^{-4t} [Using key points]

\therefore {\color{Blue} L^{-1}\left \{ \frac{3s}{\frac{1}{2}s^2-8} \right \}=3e^{4t}+3e^{-4t}}

Alternatively,

L^{-1}\left \{ \frac{3s}{\frac{1}{2}s^2-8} \right \}=L^{-1}\left \{ \frac{6s}{s^2-16} \right \}

=6L^{-1}\left \{ \frac{s}{s^2-4^2} \right \} [Using linearity of inverse laplace transform]

=6\cosh(4t)

\therefore {\color{Blue} L^{-1}\left \{ \frac{3s}{\frac{1}{2}s^2-8} \right \}=6\cosh(4t)}

Hope this helps!